OpenStudy (anonymous):

Help me with find dy/dx for the function : y=(3x+6)^1/2

hartnn (hartnn):

chain rule it is, know whats chain rule ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so did u try this ? whats your attempt ?

OpenStudy (anonymous):

i know what chain rule is, but im not sure if this is a chain rule

OpenStudy (anonymous):

take log on both sides ..and do diffrentiaton

hartnn (hartnn):

u know whats the derivative of sqrt x ?

OpenStudy (anonymous):



OpenStudy (anonymous):

hartnn no shkrina oookkayy, let me understand this first...

hartnn (hartnn):

know the derivative of x^n ?

OpenStudy (anonymous):

guys there r many ways to to do .. simple way sgoes by applying log ...

OpenStudy (anonymous):

other way?

hartnn (hartnn):

definition of simple is different for everyone ^_^

hartnn (hartnn):

we are discussing 2 methods, my chain rule method, and her log method

OpenStudy (anonymous):

the chain rule...umm, the step solution?

hartnn (hartnn):

for chain rule, you should be able to calculate the derivative of sqrt x...

hartnn (hartnn):

or x^(1/2) using the formula for x^n

OpenStudy (anonymous):

(3x+6)^1/2 = (3x+6)^-2 is it wrong?

hartnn (hartnn):

yes, its incorrect

OpenStudy (anonymous):

(3x+6)^1/2 = ??

OpenStudy (anonymous):



hartnn (hartnn):

let 3x+6 = f(x) so, \(\huge \dfrac{d}{dx}\sqrt{f(x)}=\dfrac{1}{2\sqrt{f(x)}}\dfrac{d}{dx}f(x)\) this is chain rule, where i have used the fact that d/dx (sqrt x) = 1/ 2sqrt x